OpenStudy (anonymous):

is f(x)=(3x+4)^(1/2) continuos and differentiable

OpenStudy (anonymous):

pretty much every function you know is continuous and differentiable on its domain

OpenStudy (anonymous):

\[f(x)=\sqrt{3x+4}\] has domain \[x\geq -\frac{4}{3}\] and it is continuous and differentiable on the open interval \(x>-\frac{3}{4}\) cannot include the left endpoint as it the limit doesn't exist from the left since the function is not defined for number less that \(-\frac{3}{4}\)

OpenStudy (anonymous):

so the mean value theorem would satisfyy it in the interval (0,4)?

OpenStudy (anonymous):

yes it would

OpenStudy (anonymous):

well actually "it would satisfy the mvt" not the other way around

OpenStudy (anonymous):

ok so my hw question is: "what value of c in the open interval (0,4) satisfies the mean value theorem for f(x)= (3x+4)^(1/2)

OpenStudy (anonymous):

The mean value theorem says there is a c in (0,4) such that f'(c) is the same as the average slope, which in this case is one half, by\[m_{average}=\frac{f(4)-f(0)}{4-0}=\frac{4-2}{4}=\frac{1}{2}\]Calculate f'(x), let f'(c)=1/2, and solve for c.